OpenStudy (nottim):

How do I find the vertex of a trinomial if there are no x-intercepts? How do I find the vertex of a trinomial if there are no x-intercepts? @Mathematics

hero (hero):

-b/(2a)

hero (hero):

x = -b/(2a) no matter what the trinomial is

hero (hero):

and then, after you find x, insert it into the trinomial to find y

hero (hero):

If you don't understand how to do that, then you should just drop out of math forever

hero (hero):

NotFlunk

OpenStudy (nottim):

waazaat yzw tkn bout?

OpenStudy (nottim):

why?

hero (hero):

x^2-2x+1 has no x-intercepts on the y axis, but it has a vertex of x = -b/2a = -(-2)/2(1) = 2/2 = 1 y = (1)^2-2(1)+1 = 1-2+1 = 2-2 = 0 so it has a vertex of (1,0)

hero (hero):

That's an example

OpenStudy (nottim):

For\[3x ^{2}-9x+16, I got 9, 178.\]

OpenStudy (nottim):

That's definietyle rong.

hero (hero):

3x^2-9x+16 x = -b/(2a) = 9/9 = 1 y = 3(1)^2-9(1) + 16 = 3-9+16 = 19-9 = 10 so the vertex of 3x^2-9x+16 = (1,10)

hero (hero):

Dude, that wasn't even close man. How did you get (9,178) ?

OpenStudy (nottim):

No according to hmy graph though...The vertex is at 1.5, 9.5 (somehwere, noexact position)

hero (hero):

Well, you must have graphed it wrong somewhere

OpenStudy (nottim):

unless you did a rough est kidna calculation.

hero (hero):

NO, I did an exact calculation. You can't roughly calculate a vertex

hero (hero):

However, you calculator can roughly estimate the graph of a function.

hero (hero):

I don't think you found the exact vertex of your graph man

OpenStudy (nottim):

noes. err, was your a 3 and your b 9?

OpenStudy (nottim):

-9

hero (hero):

Oh I found my mistake. Let me re-cal

OpenStudy (nottim):

its 1.5,9.25 right?

hero (hero):

x = -b/(2a) = -(-9)/2(3) = 9/6 = 3/2 Yeah but the sad thing is, you still didn't use the formula to find it.

hero (hero):

You had to graph it

hero (hero):

I feel sorry for you man

OpenStudy (nottim):

itws part of a larger assignment.

OpenStudy (nottim):

something with calcs and graphs.

OpenStudy (nottim):

thx anyways.